alternating current in circuit is given byi=inot sin2piwt then time taken by the current to rise from zero to maximum
 

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Current will be maximum for maximum value of sin will be maximum=1.
This occurs for value of angle =\dfrac{\pi}{2}
Hence this occurs for2\pi\omega t=\dfrac{\pi}{2}\implies t=\dfrac{1}{4\omega}